OpenStudy (anonymous):

Determine the perimeter of quadrilateral KLMN with vertices K(-4, 2), L(2, 2), M(8, -3), and N(-10, -6). Answer 42 units 23.8 units 32 units 33.2 units

OpenStudy (anonymous):

distance formula for KL, LM, KN, and NK

OpenStudy (anonymous):

d = \[\sqrt{(x2 - x1)^2 + (y2 - y1)^2}\]

OpenStudy (anonymous):

For perimeter, you need to know the lengths of the sides. KL=sqrt[(2+4)^2 + (2-2)^2] KL=6 LM=sqrt[(8-2)^2+(-3-2)^2] LM=sqrt(36+25) LM=sqrt61

OpenStudy (anonymous):

MN=sqrt[(-10-8)^2+(-6+3)^2] MN=sqrt(324+9) MN=3sqrt37

OpenStudy (anonymous):

KN=sqrt[(-10+4)^2+(-6-2)^2] KN=sqrt(36+64) KN=10 P = KL+LM+MN+KN=6+sqrt61+3sqrt37+10 P=16+3sqrt37+10